# How do Primordial Black Holes change the Halo Mass Function and Structure?

### v1

@@ 1 Introduction
Since the James Webb Space Telescope (JWST) deployment in 2022, groundbreaking observations have begun to illuminate the enigmatic early universe and the genesis of its first galaxies. Recent JWST observations have revealed an astonishing population of massive galaxies with stellar masses exceeding $10^{9}\,\mathrm{M}_{\odot}$ at redshifts $z\gtrsim 10$ (e.g., Finkelstein et al., 2022; Castellano et al., 2022; Naidu et al., 2022; Yan et al., 2023a; Castellano et al., 2023; Labbé et al., 2023; Yan et al., 2023b), and supermassive black holes (SMBHs) with masses $\gtrsim 10^{7}\,\mathrm{M}_{\odot}$ at $z=7-10$ (e.g., Goulding et al., 2023; Bogdán et al., 2024; Larson et al., 2023; Maiolino et al., 2023; Greene et al., 2024; Natarajan et al., 2024; Kovacs et al., 2024). While most galaxies are identified through photometric measurements subject to calibration uncertainties, a select few have been spectroscopically confirmed (i.e. Curtis-Lake et al., 2023; Arrabal Haro et al., 2023; Robertson et al., 2023). The existence of those galaxies, with higher than expected stellar mass at such high redshift (e.g., GN-z11; Tacchella et al. 2023), challenges our standard $\Lambda$CDM cosmology (Inayoshi et al., 2022; Boylan-Kolchin, 2023). To alleviate this tension, a re-calibration of the stellar mass has been proposed, such as using different stellar initial mass functions and their resulting spectral energy distributions, when interpreting the observations (e.g., Finkelstein et al., 2023; Wang et al., 2024).
A related challenge is posed by the discovery and follow-up spectroscopic confirmation of massive black holes in the pre-reionization universe (e.g., UHZ1; Goulding et al. 2023; Bogdán et al. 2024). The standard SMBH formation channel is via gas accretion onto the light black hole seeds from massive Population III (Pop III) star remnants with $\sim 100-1000\ \mathrm{M}_{\odot}$  (for reviews, see Smith & Bromm, 2019; Inayoshi et al., 2020). However, rapid growth of the light seeds to $10^{7}\,\mathrm{M}_{\odot}$ in less than $\sim 500$ Myr by $z\sim 10$ requires an accretion rate exceeding the Eddington limit (e.g., Bogdán et al., 2024). Recent observations of high-redshift active galactic nuclei (AGN) suggest that the majority may be dormant with sub-Eddington accretion rates (e.g., Jeon et al., 2023). If interspersed with episodic super-Eddington phases, however, such light-seed models may be able to explain at least some of the newly discovered high-$z$ SMBHs (Juodžbalis et al., 2024; Maiolino et al., 2024). An alternative SMBH formation channel proceeds via heavy black hole seeds from the direct collapse of a primordial gas cloud (Bromm & Loeb, 2003; Begelman et al., 2006; Lodato & Natarajan, 2006). The resulting accelerated growth (e.g., Jeon et al., 2024) may be required to explain part of the emerging JWST phenomenology (Bogdán et al., 2024; Kovacs et al., 2024).
@@ 2 PBH Halo Mass Function
For simplicity, we assume a monochromatic mass distribution for PBHs with $M_{\rm PBH}=10^{6}\,\mathrm{M}_{\odot}$. Observations constrain the abundance of PBHs as a fraction of total DM mass, denoted by $f_{\rm PBH}$, leading to maximum values allowed for a monochromatic distribution (Carr et al., 2021b). The current upper limit on $10^{6}\,\mathrm{M}_{\odot}$ PBHs is $f_{\rm PBH}\lesssim 10^{-3}$, deriving from the effect of dynamical friction (Carr & Sakellariadou, 1999) and the formation of large-scale structure (Carr & Silk, 2018). We also note that the most stringent limit arises from $\mu$-distortion of the cosmic microwave background (CMB) when assuming PBH formation from the high-density tail of a Gaussian random field (Nakama et al., 2018; Hooper et al., 2024). However, relaxing the Gaussian fluctuation assumption (Nakama et al., 2018) and considering formation mechanisms dependent on a different inflation field (i.e. Kawasaki & Murai, 2019; Carr & Kühnel, 2020) could alleviate this bound. Therefore, in this paper, we adopt $10^{-3}$ as the fiducial value, consistent with existing bounds, and vary the PBH fraction in DM within $f_{\rm PBH}\sim 10^{-2}-10^{-4}$ to study its impact on the halo mass function (HMF).
@@
Further work will be needed to fully explore this heuristic term.
@@ 3.2 PBH initial conditions
Here, $r$ is the distance of the PDM particle from its nearest PBH. With this prescription, we create a uniform sphere of PDM surrounding the PBH with an overdensity of $\Delta_{\rm vir}$ and a radius $d_{\rm nl}$, to approximate the nonlinear features. By running our gizmo simulations (see the next section), we find that the mass of the virialized halo around an isolated PBH grows linearly with $a$, as $M_{h}\sim(4\pi/3)D_{\rm PBH}(a)m_{\rm PBH}$, which is generally consistent with the spherical collapse theory (see Mo et al., 2010), where $M_{\rm h}\sim m_{\rm PBH}a/a_{\rm eq}$ (Mack et al., 2007), but the normalization is off by a factor of a few. To mitigate the difference in halo mass between the analytical predictions by Mack et al. (2007) and our simulations using the growth factor from Equ. (3), we correct the growth factor to $D_{\rm PBH}(a_{\rm ini})\sim 3a_{\rm ini}/(4\pi a_{\rm eq})$, as indicated by CORR in Table 1. Finally, we combine the two perturbation regimes to calculate the displacement field induced by PBHs as $\vec{\psi}=\vec{\psi}_{\rm seed}+\vec{\psi}_{\rm Poisson}$, capturing both ‘seed’ and ‘Poisson’ effects.
@@ 4.2 Large-Scale Structure
Figure 3: Comparing the simulated halo mass function (dashed lines) from different initial conditions after removing all subhalos at $z=10$. To provide a reference point, we also reproduce the HMF derived from the analytical formalism in Section 2 (solid lines) at the same redshift, noting that the same line color indicates identical initial conditions and PBH parameters. Left panel: Effect of varying the fraction of PBHs in the DM, specifically considering $f_{\rm PBH}=10^{-2}$ (PBH2), $10^{-3}$ (PBH3), and $10^{-4}$ (PBH4). Right panel: Effect of varying numerical implementation and parameters, for a fixed $f_{\rm PBH}=10^{-3}$. Specifically, we consider the absence of the correction of the growth factor (PBH_NS), non-linear perturbations from the nearest PBH (PBH_PS), and of both (PBH_PS_NS). Additionally, the softening parameter for PBH particles is increased in the PBH3_smooth case. In both panels, the horizontal dashed line represents the number density resolution limit set by the comoving volume of our box as $1/V_{\rm com}$.
@@
Subsequently, we count the halos within each mass bin and plot the HMF after subhalo removal at $z=10$, as shown in Figure 3. For ease of comparison, we also reproduce the analytical HMFs, as derived in Section 2. The left panel of Figure 3 illustrates the effects of varying the PBH mass abundance parameter $f_{\rm PBH}$, and the right panel displays the effects of different initial conditions or numerical implementations for the same $f_{\rm PBH}$. Here, all simulations are represented with dashed lines. The sharp cutoff in our simulated results can be attributed to the halo selection mass limit at the lower end and the box size limit at the higher end. From the left panel, we find that with higher $f_{\rm PBH}$, the abundance of halos with mass $\ll 10^{8}\mathrm{M}_{\odot}$ is greatly suppressed, in contrast to the analytical results. Comparing the simulation results in the right-hand panel, the differences are minimal, indicating that variations in the choice of numerical parameters do not greatly change the HMF. However, in the PBH3_NS implementation, we observe that the mass of the most massive halos is approximately twice as large as in the PBH3 run. This effect is likely due to the larger structure growth factor $D_{\rm PBH}$ without the correction motivated by the results in Mack et al. (2007, see Sec. 3.2), resulting in heavier PBH-seeded halos. At the lower mass end ($\ll 10^{8}\mathrm{M}_{\odot}$), the halo abundance in the PBH3_NS run is suppressed by a factor of 2 as compared to the PBH3 case. Both panels suggest that more massive halos suppress the abundance of less massive halos, a topic we will elaborate on in more detail in Section 4.4.
@@ 4.3 Small-Scale Structure
The remaining PDM and PBH host halos typically scale as $\rho\propto r^{-1}$ for the central profile. At larger distances from the halo center, both PBH host and PDM halos exhibit a steeper profile, with $\rho\propto r^{-2.2}$, largely consistent with the Navarro-Frenk-White (NFW) profile. A similar scaling is also observed when sampling PDM halos in the CDM run for comparison. We also examine the velocity dispersion versus normalized radius in the right panel of Figure 4. We find that the overall velocity dispersion is higher for PBH hosts, as the selected cases are more massive, and we note that PDM halos show a more flattened radial profile.
@@ 4.4 Nearest-Neighbor Statistics
Figure 7: Halo assembly in the early universe. In the $\Lambda$CDM case, minihalos are expected to cluster and undergo hierarchical merging to form larger mass halos (top panel). However, PBHs significantly disrupt this hierarchical merger process for minihalos formed close to PBH-seeded halos (bottom panel). With increasing PBH abundance, almost no minihalo can withstand this ingestion, and nearly all are likely to be absorbed by the nearest PBH-seeded halo. This dynamic interaction underscores the pivotal role of PBHs in altering the standard view of cosmic structure formation through their impact on hierarchical merging within the PBH-CDM setting.

### v2

@@ 1 Introduction
Since the James Webb Space Telescope (JWST) deployment in 2022, groundbreaking observations have begun to illuminate the enigmatic early universe and the genesis of its first galaxies. Recent JWST observations have revealed an astonishing population of massive galaxies with stellar masses exceeding $10^{9}\,\mathrm{M}_{\odot}$ at redshifts $z\gtrsim 10$ (e.g., Finkelstein et al., 2022; Castellano et al., 2022; Naidu et al., 2022; Yan et al., 2023a; Castellano et al., 2023; Labbé et al., 2023; Yan et al., 2023b), and supermassive black holes (SMBHs) with masses $\gtrsim 10^{7}\,\mathrm{M}_{\odot}$ at $z=7-10$ (e.g., Goulding et al., 2023; Bogdán et al., 2024; Larson et al., 2023; Maiolino et al., 2023; Greene et al., 2024; Natarajan et al., 2024; Kovács et al., 2024). While most galaxies are identified through photometric measurements subject to calibration uncertainties, a select few have been spectroscopically confirmed (i.e. Curtis-Lake et al., 2023; Arrabal Haro et al., 2023; Robertson et al., 2023). The existence of those galaxies, with higher than expected stellar mass at such high redshift (e.g., GN-z11; Tacchella et al. 2023), challenges our standard $\Lambda$CDM cosmology (Inayoshi et al., 2022; Boylan-Kolchin, 2023). To alleviate this tension, a re-calibration of the stellar mass has been proposed, such as using different stellar initial mass functions and their resulting spectral energy distributions, when interpreting the observations (e.g., Finkelstein et al., 2023; Wang et al., 2024).
A related challenge is posed by the discovery and follow-up spectroscopic confirmation of massive black holes in the pre-reionization universe (e.g., UHZ1; Goulding et al. 2023; Bogdán et al. 2024). The standard SMBH formation channel is via gas accretion onto the light black hole seeds from massive Population III (Pop III) star remnants with $\sim 100-1000\ \mathrm{M}_{\odot}$  (for reviews, see Smith & Bromm, 2019; Inayoshi et al., 2020). However, rapid growth of the light seeds to $10^{7}\,\mathrm{M}_{\odot}$ in less than $\sim 500$ Myr by $z\sim 10$ requires an accretion rate exceeding the Eddington limit (e.g., Bogdán et al., 2024). Recent observations of high-redshift active galactic nuclei (AGN) suggest that the majority may be dormant with sub-Eddington accretion rates (e.g., Jeon et al., 2023). If interspersed with episodic super-Eddington phases, however, such light-seed models may be able to explain at least some of the newly discovered high-$z$ SMBHs (Juodžbalis et al., 2024; Maiolino et al., 2024). An alternative SMBH formation channel proceeds via heavy black hole seeds from the direct collapse of a primordial gas cloud (Bromm & Loeb, 2003; Begelman et al., 2006; Lodato & Natarajan, 2006). The resulting accelerated growth (e.g., Jeon et al., 2024) may be required to explain part of the emerging JWST phenomenology (Bogdán et al., 2024; Kovács et al., 2024).
@@ 2 PBH Halo Mass Function
For simplicity, we assume a monochromatic mass distribution for PBHs with $M_{\rm PBH}=10^{6}\,\mathrm{M}_{\odot}$. Observations constrain the abundance of PBHs as a fraction of total DM mass, denoted by $f_{\rm PBH}$, leading to maximum values allowed for a monochromatic distribution (Carr et al., 2021b). The current upper limit on $10^{6}\,\mathrm{M}_{\odot}$ PBHs is $f_{\rm PBH}\lesssim 10^{-3}$, deriving from the effect of dynamical friction (Carr & Sakellariadou, 1999) and the formation of large-scale structure (Carr & Silk, 2018). We also note that the most stringent limit arises from $\mu$-distortion of the cosmic microwave background (CMB) when assuming PBH formation from the high-density tail of a Gaussian random field (Chluba et al., 2012; Nakama et al., 2018; Chluba et al., 2021; Hooper et al., 2024), and the future mission PIXIE will further refine the measurement of the CMB power spectrum on small scales (Kogut et al., 2024). On the other hand, relaxing the Gaussian fluctuation assumption (Nakama et al., 2018) and considering formation mechanisms dependent on a different inflation field (i.e. Kawasaki & Murai, 2019; Carr & Kühnel, 2020) could alleviate this bound. Therefore, in this paper, we adopt $10^{-3}$ as the fiducial value, consistent with existing bounds, and vary the PBH fraction in DM within $f_{\rm PBH}\sim 10^{-2}-10^{-4}$ to study its impact on the halo mass function (HMF).
@@
Further work will be needed to fully explore this heuristic term [^1] We have also compared the calculated power spectrum from Equ. (1) with that directly measured from the initial condition of our simulation box for both $\Lambda$CDM and PBH universes. The measurement for the $\Lambda$CDM simulation box is in good agreement with the Planck15 spectrum (Planck Collaboration et al., 2016) at $k\sim 10-10^{3}\,h\,\mathrm{Mpc}^{-1}$. For the PBH simulation boxes, a reasonably good convergence was found on small scales for all cases at around $k\gtrsim 100\,h\,\mathrm{Mpc}^{-1}$ , which corresponds to the isocurvature term $P_{\rm iso}(k)$ in Equ. (1). On larger scales ($k\lesssim 100\,h\,\mathrm{Mpc}^{-1}$), the measured power spectrum starts to deviate from the analytical form. Besides, limited by the box size, the contributions of large-scale modes $k\ll 10\,h\,\mathrm{Mpc}^{-1}$ are missing. We conclude that due to the nonlinear effects around individual PBHs, the heuristic term $P_{\rm corr}(k)$ in Equ. (4) might not accurately capture the perturbations by PBHs. To find a more accurate and general analytical form of $P_{\rm corr}(k)$, we need larger simulation boxes with higher resolution, which warrants future work..
@@ 3.2 PBH initial conditions
Here, $r$ is the distance of the PDM particle from its nearest PBH. With this prescription, we create a uniform sphere of PDM surrounding the PBH with an overdensity of $\Delta_{\rm vir}$ and a radius $d_{\rm nl}$, to approximate the nonlinear features. By running our gizmo simulations (see the next section), we find that the mass of the virialized halo around an isolated PBH grows linearly with $a$, as $M_{\rm h}\sim(a/a_{\rm ini})(4\pi/3)D_{\rm PBH}(a_{\rm ini})m_{\rm PBH}$, which is generally consistent with the spherical collapse theory (see Mo et al., 2010), where $M_{\rm h}\sim m_{\rm PBH}a/a_{\rm eq}$ (Mack et al., 2007), but the normalization is off by a factor of a few. To mitigate the difference in halo mass between the analytical predictions by Mack et al. (2007) and our simulations using the growth factor from Equ. (3), we correct the growth factor to $D_{\rm PBH}(a_{\rm ini})\sim 3a_{\rm ini}/(4\pi a_{\rm eq})$, as indicated by CORR in Table 1. Finally, we combine the two perturbation regimes to calculate the displacement field induced by PBHs as $\vec{\psi}=\vec{\psi}_{\rm seed}+\vec{\psi}_{\rm Poisson}$, capturing both ‘seed’ and ‘Poisson’ effects.
@@ 4.2 Large-Scale Structure
Figure 3: Comparing the simulated halo mass function (dashed lines) from different initial conditions after removing all subhalos at $z=10$. To provide a reference point, we also reproduce the HMF derived from the analytical formalism in Section 2 (solid lines) at the same redshift, noting that the same line color indicates identical initial conditions and PBH parameter. Left panel: Effect of varying the fraction of PBHs in the DM, specifically considering $f_{\rm PBH}=10^{-2}$ (PBH2), $10^{-3}$ (PBH3), and $10^{-4}$ (PBH4). Right panel: Effect of varying numerical implementation and parameters, for a fixed $f_{\rm PBH}=10^{-3}$. Specifically, we consider the absence of the correction of the growth factor (PBH_NS), non-linear perturbations from the nearest PBH (PBH_PS), and of both (PBH_PS_NS). Additionally, the softening parameter for PBH particles is increased in the PBH3_smooth case. In both panels, the horizontal dashed line represents the number density resolution limit set by the comoving volume of our box as $1/V_{\rm com}$.
@@
Subsequently, we count the halos within each mass bin and plot the HMF after subhalo removal at $z=10$, as shown in Figure 3555We have also checked the convergence of the HMF from simulations with the analytical form. General convergence is observed at around $\mathcal{O}(10^{8})\mathrm{M}_{\odot}$, and the simulated HMF shows some deviation at low masses. This difference can be explained by the non-linear effects induced by PBHs at small scales, which cannot be fully modelled by the PS formalism. Here, to show the effect of PBH seeding and bimodal features in halo mass distributions, we only plot the mass distribution after subhalo removal. . For ease of comparison, we also reproduce the analytical HMFs, as derived in Section 2666The analytical HMFs displayed in Figure 3 are zoomed-in versions of those shown in the right panel in Figure 1, selecting the most relevant mass scale from our simulations for better visualization purposes.. The left panel of Figure 3 illustrates the effects of varying the PBH mass abundance parameter $f_{\rm PBH}$, and the right panel displays the effects of different initial conditions or numerical implementations for the same $f_{\rm PBH}$. Here, all simulations are represented with dashed lines. The sharp cutoff in our simulated results can be attributed to the halo selection mass limit at the lower end and the box size limit at the higher end. From the left panel, we find that with higher $f_{\rm PBH}$, the abundance of halos with mass $\ll 10^{8}\mathrm{M}_{\odot}$ is greatly suppressed, in contrast to the analytical results. Comparing the simulation results in the right-hand panel, the differences are minimal, indicating that variations in the choice of numerical parameters do not greatly change the HMF. However, in the PBH3_NS implementation, we observe that the mass of the most massive halos is approximately twice as large as in the PBH3 run. This effect is likely due to the larger structure growth factor $D_{\rm PBH}$ without the correction motivated by the results in Mack et al. (2007, see Sec. 3.2), resulting in heavier PBH-seeded halos. At the lower mass end ($\ll 10^{8}\mathrm{M}_{\odot}$), the halo abundance in the PBH3_NS run is suppressed by a factor of 2 as compared to the PBH3 case. Both panels suggest that more massive halos suppress the abundance of less massive halos, a topic we will elaborate on in more detail in Section 4.4.
@@ 4.3 Small-Scale Structure
The remaining PDM halos in the PBH3 run are less resolved due to their low masses close to the resolution limit, whereas PBH host halos typically exhibit $\rho\propto r^{-1}$ scaling for the central profile. At larger distances from the halo center, both PBH host and PDM halos exhibit a steeper profile, with $\rho\propto r^{-2.2}$, largely consistent with the Navarro-Frenk-White (NFW) profile. A similar scaling is also observed when sampling PDM halos in the CDM run for comparison. We also examine the velocity dispersion versus normalized radius in the right panel of Figure 4. We find that the overall velocity dispersion is higher for PBH hosts, as the selected cases are more massive, and we note that PDM halos show a more flattened radial profile.
@@ 4.4 Nearest-Neighbor Statistics
Figure 7: Halo assembly in the early universe. In the $\Lambda$CDM case, minihalos are expected to cluster and undergo hierarchical merging to form higher-mass halos (top panel). However, PBHs significantly disrupt this hierarchical merger process for minihalos formed close to PBH-seeded halos (bottom panel). With increasing PBH abundance, almost no minihalo can withstand this ingestion, and nearly all are likely to be absorbed by the nearest PBH-seeded halo. This dynamic interaction underscores the pivotal role of PBHs in altering the standard view of cosmic structure formation through their impact on hierarchical merging within the PBH-CDM setting.